# Demo: SGCode: A Flexible Prompt-Optimizing System for Secure Generation of Code

### v1

@@ 3. SGCode System
We develop SGCode (Figure 1) as an API-accessible service and a web application for users to access our system on the web or integrate our API into their applications. The system design focuses on simplicity to optimize system costs, extensibility to customize the API, and maintainability with our modular architecture.
Figure 1. Basic Architecture of SGCode.
3.1. Back-end Services
@@
(a) Chat Interface
(b) Report Analysis Interface
@@ 3.2. Front-End Interface
The interface is built using Streamlit, an open-source Python framework for creating interactive user interfaces (UI). The interface is organized into two main screens: (1) AI Prompt UI, which facilitates user input (Figure 2(a)); and (2) Security Analysis Report screen, which displays detailed results and insights (Figure 2(b)).

### v3

@@ 3. SGCode System
We develop SGCode (Figure 1(a)) as an API-accessible service and a web application for users to access our system on the web or integrate our API into their applications. The system design focuses on simplicity to optimize system costs, extensibility to customize the API, and maintainability with our modular architecture.
@@
(a) Basic Architecture of SGCode.
(b) Chat Interface
(c) Report Analysis Interface
@@ 3.1. Front-End Interface
The interface is built using Streamlit, an open-source Python framework for creating interactive user interfaces (UI). The interface is organized into two main screens: (1) AI Prompt UI, which facilitates user input (Figure 1(b)); and (2) Security Analysis Report screen, which displays detailed results and insights (Figure 1(c)).